OpenStudy (exoexoexo):

a white die and a black die are thrown. what is the probability that the number on the white die exceeds the number on the black die?

OpenStudy (anonymous):

I am not 100% sure but I think you do this. (If both are six sided die)

OpenStudy (anonymous):

Each box is every possible outcome, with a probability of (1/6)*(1/6) = 1/36 of occurring. Each red square is when number on white is bigger. There are 15 altogether, therefore the probability of white die exceeding black die is 15*(1/36) = 15/36